OpenStudy (anonymous):

An electron and a proton both start from rest and from the same point in a uniform electric field of 319.4 N/C. How far apart are they 2.59 µs after they are released? Ignore the attraction between the electron and the proton. (Imagine the experiment performed with the proton only, and then repeat with the electron only.) Answer in units of m

OpenStudy (tcy):

\[s=\frac{ 1 }{ 2 }at ^{2}\] \[s _{T}=s _{e}+s _{p}\] \[s _{T}=\frac{ 1 }{ 2 }a _{p}t ^{2}+\frac{ 1 }{ 2 }a _{e}t ^{2}\] \[a=\frac{ F }{ m };F=\left|q\right| E\] \[s _{T}=\frac{ 1 }{ 2 }\frac{ F }{ m_{p} }t ^{2}+\frac{ 1 }{ 2 }\frac{ F }{ m_{e} }t ^{2}\]